# An All-digital 8.6-nJ/Frame 65-nm Tsetlin Machine Image Classification Accelerator

### v2

@@ 21 $\mu$W b,d
For power consumption measurements, a Joulescope JS220 precision energy analyzer was used. The power consumption during inference was measured while the accelerator was set in a test mode with repeated classifications of the full 10k sample dataset. The power consumed by the accelerator core with 1.20 V supply voltage and a clock frequency of 27.8 MHz was 1.15 mW. This corresponds to an EPC of 19.1 nJ. With a supply voltage of 0.82 V, the accelerator core’s power consumption was 0.52 mW and the corresponding EPC was 8.6 nJ. With 1.2 V core supply voltage and a clock frequency of 1 MHz, the achieved classification rate was 2.27k FPS at a power consumption of 81 $\mu$W.
@@
Table I summarizes the characteristics and performance of the accelerator. In a real application, the ConvCoTM design would typically be included as a peripheral module to a microcontroller in an SoC. Therefore, we considered the power consumption of the accelerator core as the most interesting power parameter. In inference mode, the digital 3.3 V IOs, which were required for interfacing to the FPGA board, had a power consumption of 0.76 mW, at a clock frequency of 27.8 MHz.
@@
All power and energy measurement results of the ASIC are average values, and were performed with clock-gating and CSRF enabled (as shown in Fig. 2). When operating at 27.8 MHz, clock-gating reduced the power consumption by approximately 60%, while the CSRF alone provided less than 1% power reduction.
VI Discussions
We consider the ConvCoTM accelerator with 128 clauses as a small configuration. In light of this, we are satisfied with the obtained test accuracy which is comparable with other low-complexity and ULP solutions. Table II shows the performance and properties of state-of-the-art ULP IC implementations for classification of MNIST and CIFAR-10 datasets. Inspired by the TinyML [39] benchmarks for embedded systems, we specify the classification rate, the peak power consumption and the EPC. These are system parameters which are independent of the accelerator model and architecture [40]. However, when used for comparing solutions, one must be careful to specify the ML task performed. The CIFAR-10 dataset is significantly more challenging than MNIST. Therefore, accelerators that operate on CIFAR-10, have lower classification rates and higher EPC compared to accelerators that operate on MNIST.
TABLE II: Comparison of the ConvCoTM IC accelerator with previous works.
This work TCAS-I´25 Zhao [20] TCAS-II’23 Yejun [21] JSSC’23       Yang [9] TCAS-I’20 Mauro [6] JSSC’21       Knag [7] TCAS-I’20 Bankman [5]
Technology 65 nm CMOS 28 nm CMOS 65 nm CMOS 40 nm CMOS 22 nm FD-SOI 10 nm FinFET 28 nm CMOS
Active area 2.7 mm2 0.261 mm2 0.57 mm2 0.98 mm2 2.3 mm2 0.39 mm2 4.6 mm2
Algorithm ConvCoTM CNN SNN Ternary CNN BNN BNN BNN
Design type Digital Analog, time domain Neuromorphic mixed-signal IMC, mixed-signal Digital (SoC) Digital IMC, mixed-signal
Image dataset MNIST, FMNIST, KMNIST MNIST MNIST MNIST CIFAR-10 CIFAR-10 CIFAR-10
Measured test accuracy 97.42%, 84.54%, 82.55% 97.9% 95.35% 97.1% 99% of nominal accuracy 86% 86%
Classifications per second 60.3k a           2.27k b 3508 233k (1.2 V)           40k (0.7 V) 549 15.4 Not stated 237
Latency e 25.4 $\mu$s a           0.66 ms b Not stated Not stated Not stated Not stated Not stated Not stated
Inference power 1.15 mW a,c 0.52 mW a,d 81 $\mu$b,c 21 $\mu$b,d 11.6 $\mu$W 9.37 mW (1.2 V) 0.517 mW (0.7 V) 96 $\mu$W 674 $\mu$f 5.6 mW (13 MHz) 607 mW (622 MHz) 0.9 mW
Energy per classification (EPC) 19.1 nJ a,c           8.6 nJ a,d           35.3 nJ b,c           9.6 nJ b,d 3.32 nJ 40.17 nJ (1.2 V) 12.92 nJ (0.7 V) 0.18 $\mu$J 43.8 $\mu$J Not stated 3.8 $\mu$J
a 27.8 MHz clock frequency. b 1.0 MHz clock frequency. c The accelerator core supply voltage (vdd) is 1.20 V. d vdd=0.82 V.e For processing of a single image sample, including data transfer. f Full SoC.
Our ConvCoTM accelerator ASIC achieves an EPC of 8.6 nJ at 0.82 V operating voltage. This is the second lowest EPC reported, for a manufactured ASIC that operates on MNIST. As the ConvCoTM ASIC is fully digital, the classification rate and thereby also the peak power consumption, can be scaled by adjusting the clock frequency. At a clock frequency of 1 MHz, the power consumption is 81 $\mu$W.
VII Scaling
@@
The ConvCoTM accelerator’s architecture described in this paper, is highly suited for scaling, and can be modified to operate on larger and more complex images. There is active research on how to apply the TM algorithm for larger and more complex images. The TM version that currently achieves the best test accuracy on CIFAR-10 is TM Composites [17, 18]. Here, several TM Specialists are applied to an image sample. Examples of specializations are different booleanization techniques and convolution window sizes. After an image has been processed, the class sums from each TM Specialist are normalized, followed by summation per class. Finally, an argmax operator is applied on the composite class sums for output prediction.
@@
We assume the same 65 nm CMOS technology is used as for the current ConvCoTM ASIC. The chip area for the CIFAR-10 accelerator is estimated by multiplying the chip area of the current ASIC (2.7 mm2) with the ratio, $R$, given by the model size of one TM Specialist (32.5 kilobyte) divided by the model size of the ASIC in this paper (5.6 kilobyte). We therefore get $R=32.5/5.6\approx 5.8$. This is a reasonable assumption because the model storage in registers (for the active TM Specialist) and the clause logic, dominate the chip area. We further assume 2 mm2 in additional chip area for booleanization logic, additional adders for class sum generation, and RAM modules for model storage.
@@
An estimate of the power consumption is found by multiplying the power consumption of the current ASIC’s accelerator core, operating at 27.8 MHz (1.15 mW), with the factor $R$. During model loading and booleanization, we assume power consumption similar to that during inference. An estimate of the energy per classification is then $R\times 1.15$ mW/(3440 FPS) $\approx 1.9~{}\mu$J/frame.
@@
per TM Specialist kilobyte literals. The address
@@
Model size: Weights 12.5k 10 classes, 1000 clauses
per TM Specialist kilobyte and 10-bit weights.
@@
Accelerator core 6.7 mW Inference mode,
power consumption 27.8 MHz system clock.
Accelerator core 1.9 $\mu$J Inference mode,
energy per 27.8 MHz system clock.
classification.

### v3

@@ 21 $\mu$W b,d
For power consumption measurements, a Joulescope JS220 precision energy analyzer was used. The power consumption during inference was measured while the accelerator was set in a test mode with repeated classifications of the full 10k sample dataset. The power consumed by the accelerator core with 1.20 V supply voltage and a clock frequency of 27.8 MHz was 1.15 mW. This corresponds to an EPC of 19.1 nJ. With a supply voltage of 0.82 V, the accelerator core’s power consumption was 0.52 mW and the corresponding EPC was 8.6 nJ.
@@
Table II summarizes the characteristics and performance of the accelerator. In a real application, the ConvCoTM design would typically be included as a peripheral module to a microcontroller in an SoC, and the digital IO interface would have been optimized. Therefore, we considered the power consumption of the accelerator core as the most interesting power parameter. In inference mode, the digital 3.3 V IOs, which were required for interfacing to the FPGA board, had a power consumption of 0.76 mW, at a clock frequency of 27.8 MHz.
@@
All reported power and energy measurement results of the ASIC are average values, and were performed with clock-gating and CSRF enabled (as shown in Fig. 4). When operating at 27.8 MHz, clock-gating reduced the power consumption by approximately 60%, while the CSRF alone provided less than 1% power reduction.
VI Possible Extensions
In this section, we describe a technique which limits the number of literals that are input to each clause [42]. This can enable a significant reduction in model size and chip area, and we provide estimates for a 28 nm CMOS implementation of the ConvCoTM accelerator. We also discuss how the accelerator could be modified to support on-device training. Finally, we describe concepts and estimates for a scaled-up solution intended for operation on the CIFAR-10 dataset.
VI-A Estimated Accelerator Performance if Implemented in 28 nm CMOS
In the current ConvCoTM ASIC accelerator, each clause takes all the literals as input, ref. Fig. 4. This makes it necessary to store and use also the exclude TA actions of the model. However, for a given clause, the literals that are excluded do not contribute to the the clause value generation and therefore represent an overhead of the model.
TM models are in general highly sparse. For example, in the MNIST model used for the ConvCoTM accelerator, 88% of the TA actions are exclude. There is active research on how to utilize the TM model sparsity to achieve more compact HW and SW solutions. One option is to use a TM training setting that specifies a maximum number of literals per clause [42], with only negligible loss of accuracy. This enables a significant reduction of the model size and the associated chip area for storage and processing of the TA actions.
As an example, we here consider the current ASIC and assume that each clause only takes 10 literals as input. To select a single literal, one would need a 272-to-1 multiplexer, and a 9-bit literal address (multiplexer selection word). That is, one would need $10\times 9=90$ bits per clause, for storing which literals are included. The principle for the clause logic could be as shown in Fig. 11. In this case, an area reduction of $(272-90)/272\approx 67$% of the TA action part of the model is achieved. In the current ASIC, the TA action part of the model, and the associated clause logic, constitute about 70% of the accelerator core’s chip area. Using the principle in Fig. 11, one could achieve a total reduction of the core area of approximately 47%.
Figure 11: Circuitry for a single clause, where only 10 literals are used as input. $s_{0}$ to $s_{9}$ are multiplexer selection signals given by the model.
A reduction in chip area and EPC can be achieved by implementing the ConvCoTM accelerator in a more advanced technology. We here estimate the chip area, power consumption and EPC if a 28 nm CMOS technology was used. With Dennard scaling [43], we roughly estimate a general reduction factor of the active chip area of $(28/65)^{2}$. In addition, we assume the number of literals per clause is limited to 10 as described above. Thus gives an estimated chip area of 0.27 mm2. At an operating frequency of 27.8 MHz (used here for comparison as this was the maximum clock rate tested for the manufactured ConvCoTM ASIC), and 0.7 V supply voltage, we roughly estimate a 50% reduction in power consumption compared to the 65 nm chip operating at 0.82 V. This results in an EPC of 4.3 nJ which would be close to what is achieved for the SNN accelerator in [20].
VI-B Extension of the ASIC Design to Support Training
Based on the ConvCoTM training concept and circuit modules used in the FPGA accelerator in [12], we here describe how the ASIC could be modified to support on-device training.
During the patch generation phase, one has to perform random patch selection per clause. In [12], this is performed based on the reservoir sampling algorithm [44]. One would need one or more RAM modules for storing all of the 361 generated patches, each consisting of 136 feature bits. In addition, a register of 9 bits is required per clause to store the address of the randomly selected patch.
For each clause, 272 TAs would be needed, i.e., one TA for each literal. We here assume that 8-bit TAs are used, based on the structure shown in Fig. 1. In [12], all TAs belonging to a single clause are updated simultaneously. Therefore, the TAs can be implemented by parallel operation of 34 single-port RAM modules, each with a word width of 64 bits, supporting 8 TAs. The number of rows in the TA RAMs would be 128, corresponding to the number of clauses in the ConvCoTM configuration used.
The model registers for the weights and the TA action signals, are already included in the ASIC, but to support training, register update logic would be required.
For random number generation during the reservoir sampling and during the clause updating phase, linear feedback shift-registers (LFSRs) of minimum 16 bits, are needed [12]. During the clause updating, one LFSR is used for the random decision of whether a clause’s TAs should receive update, and 272 LFSRs are required for the simultaneous updating of all TAs per clause.
The additional circuit modules required for implementing the training modules is moderate, and we estimate it to approximately 1 mm2. For inference, the power consumption should not increase much, provided clock-gating is applied for the specific training modules, and RAM modules with very low leakage current are used for the TAs. Reading from and writing to the RAMs consume energy, but these operations would only be performed during training. The accelerator in [12] operated at 50 MHz and achieved a training throughput of approximately 40k samples per second. For a modified ASIC that supports training, and operates at 27.8 MHz, we estimate a training throughput of 22.2k samples per second.
VI-C Scaled-Up Solution
@@
The ConvCoTM accelerator’s architecture described in this paper, is highly suited for scaling, and can be modified to operate on larger and more complex images. There is active research on how to apply the TM algorithm on larger and more complex images. The TM version that currently achieves the best test accuracy on CIFAR-10 is TM Composites [17, 18]. Here, several TM Specialists are applied to an image sample. Examples of specializations are different booleanization techniques and convolution window sizes. After an image has been processed, the class sums from each TM Specialist are normalized, followed by summation per class. Finally, an argmax operator is applied on the composite class sums for output prediction.
@@
We first assume the same 65 nm CMOS technology is used as for the current ConvCoTM ASIC. The chip area for the CIFAR-10 accelerator is estimated by multiplying the chip area of the current ASIC (2.7 mm2) with the ratio, $R$, given by the model size of one TM Specialist (32.5 kilobytes) divided by the model size of the ASIC in this paper (5.6 kilobytes). We therefore get $R=32.5/5.6\approx 5.8$. This is a reasonable assumption because the model storage in registers (for the active TM Specialist) and the clause logic, dominate the chip area. We further assume 2 mm2 in additional chip area for booleanization logic, additional adders for class sum generation, and RAM modules for model storage.
@@
An estimate of the power consumption is found by multiplying the power consumption of the current ASIC’s accelerator core, operating at 27.8 MHz and 0.82 V, with the factor $R$. During model loading and booleanization, we assume power consumption similar to that during inference. An estimate of the EPC is then $R\times{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}0.52}$ mW/(3440 FPS) $\approx{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}0.9}~{}\mu$J/frame.
For a 28 nm CMOS implementation of this solution, we roughly estimate an accelerator core area of 3.3 mm2, based on Dennard scaling. Assuming 0.7 V supply voltage, a coarse estimate of the EPC is 0.45 $\mu$J. The operation frequency could be increased for higher throughput, and this could enable further reduction of the EPC.
@@
per TM Specialist kilobytes literals. The address
@@
Model size: Weights 12.5 10 classes, 1000 clauses
per TM Specialist kilobytes and 10-bit weights.
@@
3.3 mm2 28 nm CMOS.
Accelerator core 3.0 mW 27.8 MHz, 0.82 V, 65 nm.
power consumption 1.5 mW 27.8 MHz, 0.7 V, 28 nm.
Accelerator core 0.9 $\mu$J 27.8 MHz, 0.82 V, 65 nm.
energy per 0.45 $\mu$J 27.8 MHz, 0.7 V, 28 nm.
classification (EPC).
VII Discussions
TABLE IV: Comparison of the ConvCoTM IC accelerator with previous works.
This work Envisaged design. This work scaled to 28 nm. TCAS-I´25 Zhao [20] TCAS-II’23 Yejun [21] JSSC’23       Yang [9]
Technology 65 nm CMOS 28 nm CMOS 28 nm CMOS 65 nm CMOS 40 nm CMOS
Active area 2.7 mm2 0.27 mm2 0.261 mm2 0.57 mm2 0.98 mm2
Algorithm ConvCoTM ConvCoTM CNN SNN Ternary CNN
Design type Digital Digital Analog, time domain Neuromorphic mixed-signal IMC, mixed-signal
Image dataset MNIST, FMNIST, KMNIST MNIST, FMNIST, KMNIST MNIST MNIST MNIST
Measured test accuracy 97.42%, 84.54%, 82.55% 97.42%, 84.54%, 82.55% 97.9% 95.35% 97.1%
Classifications per second 60.3 k a           2.27 k b 60.3 k a 3508 233 k (1.2 V)           40 k (0.7 V) 549
Latency e 25.4 $\mu$s a           0.66 ms b 25.4 $\mu$s a Not stated Not stated Not stated
Inference power 1.15 mW a,c 0.52 mW a,d 81 $\mu$b,c 21 $\mu$b,d 0.26 mW a,f 11.6 $\mu$W 9.37 mW (1.2 V) 0.517 mW (0.7 V) 96 $\mu$W
Energy per classification (EPC) 19.1 nJ a,c           8.6 nJ a,d           35.3 nJ b,c           9.6 nJ b,d 4.3 nJ a,f 3.32 nJ 40.17 nJ (1.2 V) 12.92 nJ (0.7 V) 0.18 $\mu$J
a 27.8 MHz clock frequency. b 1.0 MHz clock frequency. c The accelerator core supply voltage (vdd) is 1.20 V. d vdd=0.82 V. e For processing of a single image sample, including data transfer. f vdd=0.7 V.
TABLE V: Comparison of the envisaged scaled-up ConvCoTM IC accelerator with previous works that operate on CIFAR-10.
Envisaged design from Subsection VI-C TCAS-I’20 Mauro [6] JSSC’21       Knag [7] TCAS-I’20 Bankman [5] TCAS-I’25 Park [26] JSSC’25 Yoshioka [27]
Technology 65 nm CMOS 28 nm CMOS 22 nm FD-SOI 10 nm FinFET 28 nm CMOS 65 nm CMOS 65 nm CMOS
Active area 17.7 mm2 (65 nm) 3.3 mm2 (28 nm) 2.3 mm2 0.39 mm2 4.6 mm2 0.17 mm2 0.48 mm2
Algorithm ConvCoTM BNN (SW defined) BNN BNN SNN (spiking VGG-16) CNN / Transformer
Design type Digital Digital (SoC) Digital IMC, mixed-signal Analog, with time-domain IMC Analog, IMC
Image dataset CIFAR-10 CIFAR-10 CIFAR-10 CIFAR-10 CIFAR-10 CIFAR-10
Test accuracy 79% 99% of nominal accuracy 86% 86% 91.13% 91.7% j           95.8% k
Classifications per second 3440 a 15.4 Not stated 237 Not stated Not stated
Latency b 0.3 ms a Not stated Not stated Not stated Not stated Not stated
Inference power 3.0 mW a,c (65 nm) 1.5 mW a,d (28 nm) 674 $\mu$e 5.6 mW l 607 mW m 0.9 mW 0.55 mW i Not stated
TOPS/W Not stated f 13 e,n 617 l         269 m 532 701.7 g         488.7 h 4094 j         818 k
Energy per classification (EPC) 0.9 $\mu$a,c (65 nm) 0.45 $\mu$a,d (28 nm) 43.8 $\mu$J Not stated 3.8 $\mu$J Not stated Not stated
a 27.8 MHz clock frequency. b For processing of a single image sample, including data transfer. c vdd=0.82 V. d vdd=0.7 V. e For the complete SoC. f Not stated due to the very different architecture compared to DNNs and SNNs. g Macro. h System. i 50 MHz, 0.75 V. j CNN. k Transformer. l 13 MHz, 0.37 V. m 622 MHz, 0.75 V. n 0.5 V.
We consider the ConvCoTM accelerator with 128 clauses as a small configuration. In light of this, we are satisfied with the obtained test accuracy which is comparable with other low-complexity and ULP solutions. Table IV shows the performance and properties of state-of-the-art ULP IC implementations for classification of MNIST. There, we have also included the performance of the envisaged 28 nm CMOS implementation of the ConvCoTM accelerator, described in Subsection VI-A. In Table V the envisaged accelerator design from Subsection VI-C is compared with other ULP accelerators that operate on the CIFAR-10 dataset.
Inspired by the TinyML [46] benchmarks for embedded systems, we specify the classification rate, the peak power consumption and the EPC. These are system parameters which are independent of the accelerator model and architecture [47]. However, when used for comparing solutions, one must be careful to specify the ML task performed. The CIFAR-10 dataset is significantly more challenging than MNIST. Therefore, accelerators that operate on CIFAR-10, have lower classification rates and higher EPC compared to accelerators that operate on MNIST.
Our ConvCoTM accelerator ASIC achieves an EPC of 8.6 nJ at 0.82 V operating voltage. This is the second lowest EPC reported, for a manufactured ASIC that operates on MNIST. As the ConvCoTM ASIC is fully digital, the classification rate and thereby also the peak power consumption, can be scaled by adjusting the clock frequency. At a clock frequency of 1 MHz and a supply voltage of 0.82 V, the achieved classification rate was 2.27 k FPS at a power consumption of 21 $\mu$W.
The estimated EPC of the envisaged design described in Subsection VI-C, for operation on CIFAR-10 images, is very low. However, the test accuracy achieved with the TM algorithm for multi-channel images, is still not at the same level as for CNNs, BNNs and SNNs. Further research is required to obtain improvement on this performance parameter. The accelerators in [7, 26, 27] achieve very high performance in terms of TOPS/W. However, as the classification rates are not specified, one cannot easily calculate the EPC numbers for these solutions.
An overview of select TM-based HW solutions, including this work, is shown in Table VI. The solutions span from all-digital to mixed-signal IMC. There are two silicon-proven ASICs and many FPGA solutions. Several TM concepts have been evaluated based on simulations.
TABLE VI: Overview of TM-based HW solutions.
This work Phil. Trans. A’20 Wheeldon [11] TCAS-I’25 Mao [31] TCAS-I’25 Tunheim [12] ISTM’23 Sahu [29] MICPRO’23 Tunheim [28] ISLPED’23 Ghazal [35] Phil. Trans. A’25 Ghazal [36]
Technology and design status ASIC, 65 nm CMOS, silicon proven ASIC, 65 nm CMOS, silicon-proven FPGA FPGA FPGA FPGA ASIC, simulation ASIC, simulation
Algorithm ConvCoTM Vanilla TM Vanilla TM, CoTM ConvCoTM Vanilla TM CTM Vanilla TM CoTM
Operation Inference Inference Training and inference Training and inference Inference Training and Inference Inference Inference
Design type Digital Digital Digital Digital Digital Digital IMC (ReRAM), mixed-signal IMC (Y-flash), mixed-signal
Dataset MNIST, FMNIST, KMNIST Binary IRIS MNIST, FMNIST, KMNIST MNIST, FMNIST, KMNIST MNIST 2D Noisy       XOR c MNIST, FMNIST, KMNIST, KWS-6 d MNIST
Test accuracy 97.42%, 84.54%, 82.55% 97.0% 97.74%, 86.38%, 83.11% 97.6%,   84.1%,   82.8% 97.71% 99.9% 96.48%, 87.67%, 88.6%, 87.1% 96.3%
Classifications per second 60.3 k a Not stated 22.4 k 134 k Not stated f 4.4 M Not stated Not stated
Inference power 0.52 mW a,b Not stated 1.65 W 1.8 W Not stated 2.529 W Not stated Not stated
Energy per classification 8.6 nJ a,b Not stated e 73.6 $\mu$J 13.3 $\mu$J Not stated 0.6 $\mu$J 13.9 nJ, 23.66 nJ, 26.47 nJ, 5.91 nJ Not stated g
a 27.8 MHz clock frequency. b vdd=0.82 V c $4\times 4$ booleanized images. d Keyword spotting dataset with six words. e 62.7 TOP/J. f The number of clock cycles required per classification is stated, for different number of clauses and with/without literal budgeting. g 24.56 TOPS/W.